# Physics of droplet regulation in biological cells

### v1

@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [187] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecule [188, 189, 190, 54], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [191] and porous networks [192]. The complex compositional features resulting from this complex mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different droplets within a cell.
@@ 3.1 Complex molecular interactions
Condensates contain a huge variety of components, including water, ions, small peptides, proteins and flexible polymers such as RNA and DNA [59, 22]. Since it is challenging to directly study interactions between these molecules in their native environment, much experimental work has thus far relied on in vitro reconstitution [193, 194, 126, 37]. Here, it is crucial to exploit label-free techniques [195, 196] since fluorescent labels typically strongly affect droplet condensation [196, 197]. One key insight from these approaches is that often a few scaffold components control the formation of condensates, while client molecules partition into them [198, 199, 200, 201, 202]. However, this dichotomy is likely an oversimplification and it will be interesting to more rigorously quantify the contributions of individual components, e.g., using the recently developed dominance metric [203], which can uncover basic interaction motifs between components [204].
@@
It is clear that biomolecular phase separation involves many different kinds of interactions between molecules [205]. To distill the essential features of these interactions, it is often helpful to coarse-grain particular microscopic models [206, 207]. For instance, many biomolecules can be conceptualised as associative polymers [208, 209, 210, 211, 81, 205] and described by sticker-and-spacer models [208, 212, 82, 22], where flexible spacers link stickers with specific interactions. However, there are also many other approaches to modelling the complex interplay of biomolecules that leads to phase separation [22, 85, 59, 213]. Such models allow us to elucidate how particular molecular features, like the amino acid sequence of proteins, affect condensate properties [214, 215], and can explain why dilute phases often exhibit surprisingly large clusters [216, 217].
@@
Interactions between biomolecules can be roughly divided into weak, transient interactions and strong, persistent bonds. While weak multivalent interactions are generally responsible for phase separation into liquid-like phases, strong interactions tend to form long-lived assemblies, which can persist as nano-domains inside condensates [218]. Moreover, chemical modifications of biomolecules can affect their interactions. For instance, proteins are changed by post-translational modifications [187], which is implicated in condensate regulation [219, 53, 220, 35]. Such changes can induce multimerisation, which can in turn enable phase separation. To understand this phenomenon, we imagine a molecule whose weak interactions are insufficient for phase separation. A multimer of such molecules, however, can interact with many other multimers, which effectively increases the molecular volume $\Omega$ while maintaining the interaction strength $\chi$ per unit volume [164]. This process essentially reduces the entropic cost of concentrating molecules, promoting phase separation. This behaviour has been used to induce artificial condensate in cells [13, 221]. More generally, it is likely that combinations of weak and strong interactions enable fine-tuning of condensate properties [222, 223].
@@ 3.3 Complex interfaces
Interfaces between droplets and the surrounding fluid are characterised by a sharp change in composition over a finite spatial width. Typically, a rapid and continuous exchange of components between droplets and surrounding fluid occurs across the interface, even at equilibrium. Since droplets in biological cells generally comprise long, complex molecules, however, this exchange may be limited by an interface resistance. Such a resistance could result from potential barriers, diffusivity reduction [256], saturation of binding sites [257], or electrostatic effects [258].
Many proteins molecules are amphiphilic, and can thus act as surfactants, preferentially associating to the interface [259, 260] to form Pickering emulsions. In general, surfactants stabilise droplet interfaces, resisting both dissolution and coalescence. The formation of complex interfaces via preferential protein association can therefore improve structural integrity of condensates and assist in size regulation [261], as well as leading to complex droplet morphologies [262]. Note that although surfactants usually result in kinetic droplet stabilisation, some complex particles may also result in thermodynamic stabilisation [263, 264].
@@ 3.4 Electrostatic effects
Charge is another important property of biomolecules that can be used to regulate condensates [269]. On the one hand, charge patterns on protein surfaces can mediate specific, short-ranged interactions [270]. On the other hand, a net charge induces electrostatic interactions that can potentially range over longer distances. Indeed, condensates can be charged [271, 272], which affects interfaces [273, 274, 258, 275], slows down coarsening [276, 277], limits droplet size [154, 278, 279], influences droplet–environment interactions [280, 281], and can potentially lead to division [282, 283]. The dynamics of charged species can be described within the simple picture of electrostatics. Denoting the charge per unit volume of species $i$ as $q_{i}$, the total charge density reads $\varrho(\boldsymbol{r})=\sum_{i}q_{i}\phi_{i}$, and induces an electrostatic potential $\varphi(\boldsymbol{r})$, which satisfies the Poisson equation
@@
where $\varepsilon$ is the electric permittivity. The potential $\varphi$ gives rise to an electric field $\boldsymbol{E}=-\nabla\varphi$ and modifies the overall free energy,
$F_{\mathrm{e}}[\Phi]=F_{\mathrm{bulk}}[\Phi]+\frac{1}{2}\int\varphi(\boldsymbol{r})\varrho(\boldsymbol{r})\text{d}V\;,$ (3.5)
where the first term is given by equation (2.6), while the second term captures electrostatic effects. The associated chemical potential, $\mu_{\mathrm{e}}=\nu\delta F_{\mathrm{e}}/\delta\phi_{i}$, gains a term proportional to $\varphi$, $\mu_{\mathrm{e}}=\mu+q_{i}\varphi$, and is thus called the electrochemical potential. This potential illustrates how positively charged molecules move down gradients in the electrostatic potential $\varphi$. In particular, small mobile ions will tend to distribute so that the system becomes charge neutral, inhibiting long-ranged electrostatic effects. In the simplest case, the ions can be described by Debye-Hückel theory, which accounts for their translational entropies and electrostatic interactions [284, 285]. The main result is that electrostatic effects do not extend significantly beyond the Debye length $\lambda_{\mathrm{D}}=[\varepsilon k_{\mathrm{B}}T/(\sum_{i}c_{i}q_{i}^{2})]^{1/2}$, where $c_{i}$ is the mean number concentration of the ion species $i$ with charge density $q_{i}$. Note that ion concentrations can depend on the phase [269], so that electrostatic interactions may be more long-ranged than expected from the average salt concentration [279].
Over short distances, electrostatic effects mediate attraction between oppositely charged molecules, which can lead to phase separation from a neutral solvent via complex coacervation [286, 25, 287, 288]. Coacervates have been studied for about a century [289] and have been implicated in the origin of life [290, 291]. In the simplest case, the attraction of the oppositely charged molecules can be captured by an additional term in the local free energy $f$ given by equation (2.1). For instance, in $d=3$ dimensions Voorn–Overbeek theory leads to the free energy density [292, 293]
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of the electrostatic interaction. This energy can be written in terms of the Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$, which describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [294]. Since the permittivity of cytosol is comparable to water [295], a typical Bjerrum length is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{\upmu m}$. We thus have $\alpha\sim 10^{3}\ldots 10^{4}$ for typical molecular volumes of $\nu_{0}\sim 10^{4}\ldots 10^{6}\,\mathrm{nm^{3}}$, indicating that electrostatic effects can lower the free energy significantly, even if the net charge of the involved biomolecules is small.
@@ 3.5 Multiphase coexistence
Cells often comprise many different types of condensates, consistent with Gibbs phase rule, which limits the number of different phases to be smaller or equal to the number of different components; see section 2.1.1. The number of coexisting phases and their compositions can generally be determined by minimising the free energy of the entire system, which leads to the coexistence conditions given by equations (2.5). However, even for the simple Flory–Huggins theory given by equation (2.1), these relations are difficult to solve since they are non-linear and depend on all interaction parameters $\chi_{ij}$ and compositions $\bar{\phi}_{i}$. It is thus not surprising that the resulting phase diagrams, showing the equilibrium phases as a function of the systems’ composition, can be complex [103]. Even in the simplest case of two interacting solutes, the topology of the phase diagram depends on the interaction strength; see Figure 3.1. This multi-component, multi-phase problem has been tackled by linear stability analysis of the homogeneous state [296, 297, 298, 118, 299], grand-canonical simulations [300, 301], continuation methods [302], convexification of the free energy density [103], and machine learning methods [303].
@@ 3.5.1  Random mixtures
where $\alpha>0$ is a shape parameter. The equimolar case is recovered in the limit $\alpha\rightarrow\infty$; $\alpha=1$ samples all compositions uniformly; and $\alpha<1$ result in mixtures with a few dominating species. Taken together, this protocol enables studying the average behaviour of a multi-component liquid with random interactions.
@@ 3.5.2  Instability of homogeneous states
Phase separation of a multi-component mixture must occur when the homogeneous state is unstable, which is the case when the Hessian $H_{ij}=\partial^{2}f/\partial\phi_{i}\partial\phi_{k}$ exhibits at least one negative eigenvalue. If there are multiple negative eigenvalues, the eigenvectors generally point in different directions, suggesting that the system segregates into various phases. The number of negative eigenvalues, $N_{\mathrm{u}}$, can be predicted for the Flory–Huggins model with
$H_{ij}$ $=\frac{k_{\mathrm{B}}T}{\nu_{0}}\left(\frac{\delta_{ij}}{\Omega_{i}\phi_{i}}+\frac{1}{\Omega_{0}\phi_{0}}+\chi_{ij}\right)\;;$ (3.8)
which follows from equation (2.1). Using the random interactions and compositions introduced in the previous subsection, one can use random matrix theory [304] to predict $N_{\mathrm{u}}$ [296, 118]. For many components and uniformly distributed compositions ($\alpha=1$), the average value of $N_{\mathrm{u}}$ is approximately [305]
@@ 3.5.3  Coexisting equilibrium phases
The scaling predicted for the number of unstable modes and the phase count is consistent with the results for binary phase separation discussed in section 2.1.3: equation (2.18) implies that the fraction above which the homogeneous state becomes unstable scales inversely with $\chi$, while equation (2.21) shows that the minimal fraction for forming a droplet also scales exponentially with $\chi$. The same scalings are captured by equations (3.9) and (3.10), suggesting that this is a universal feature of phase-separating systems. This scaling also implies that a stability analysis of the homogeneous state does not carry reliable information about the phase behaviour of multi-component liquids.
@@ 3.5.4  Towards realistic systems
The scaling result given by equation (3.10) predicts that the number ${N_{\mathrm{p}}}$ of coexisting phases scales with the component count ${N_{\mathrm{c}}}$ in random mixtures. Yet, such scaling is not observed in cells, where ${N_{\mathrm{c}}}\gtrsim 10^{5}$ [187], but there are likely ${N_{\mathrm{p}}}\lesssim 10^{2}$ different kinds of condensate. This discrepancy suggests that the simple random approximations are insufficient to understand phase separation in cells. This is not surprising since the cell’s composition is clearly not random, but homeostatically controlled. Moreover, interactions between biomolecules are not random, but likely structured [306, 307, 308, 309] and also under cellular control [97, 310, 311, 312]. Interactions are also typically complex [313], likely involving higher-order interactions [314] and polymer dynamics [315, 211, 209]. Note also that the number ${N_{\mathrm{p}}}$ of coexisting phases is only the most basic descripter of multi-component phase separation, while the compositions of phases are also crucial for cellular function since they give rise to the droplets’ identities. Along those lines, it will also be interesting to understand what determines the interfacial parameters $\kappa_{ij}$ in equation (2.6), which are crucial for predicting surface tensions and thus morphologies [316, 103, 317]. In the simplest case, surface tension and related quantities might simply scale as a power law of the distance to a critical point [318], but whether this holds generally is unclear. Besides surface tension, a droplets’ morphology is also shaped by other objects in its surrounding, as discussed in the next section.
@@ 4.1.1  Partial, full and pre- wetting
Figure 4.1: Different phenomena resulting from droplet–structure interactions. (A) Dense phase partially wets the external boundary, forming a compact droplet whose interface meets the boundary at a contact angle $\vartheta$. The area of the interfaces between the droplet and the dilute phase, and the droplet and the boundary are denoted by $A$ and $S$, respectively. (B) Dense phase completely wets the boundary. (C) In an undersaturated system, droplet material pre-wets the boundary, forming a layer of thickness $h$. (D) Droplet volume fraction $\phi$ as a function of distance from the boundary. Note that the gradient of $\phi$ at $z=0$ is the same for all three cases. (E) Phase diagram showing the pre-, full and partial wetting regimes. The typical size of the pre-wetting regime has been exaggerated for clarity. (F) Droplet–droplet interactions between two immiscible phases result in either droplet separation, engulfment, or partial wetting (clockwise from top left). Panels (A) to (E) inspired by figures in [320] and panel (F) inspired by figure in [13].
@@
In a supersaturated system ($\bar{\phi}>\phi^{(0)}_{\mathrm{out}}$), $\Delta\!f>0$, and so macroscopic droplets are able to form regardless of the presence of external boundaries. Nevertheless, surface interactions still have a significant impact on droplet regulation, both in terms of droplet nucleation (see section 4.1.4 below) and droplet morphology. In particular, if $\gamma_{\mathrm{in}}<\gamma_{\mathrm{out}}$ and $\gamma<\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}$, the dense phase will completely ‘wet’ the surface, forming a macroscopic layer between the surface and the dilute phase. This configuration is known as perfect, or full, wetting. For systems in which $\gamma>|\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}|$, the dense phase will instead partially wet the surface to form spatially compact droplets. The crossover between these two cases is known as the wetting transition [321]. The different phenomenologies of pre-, partial, and full wetting are highlighted in figure 4.1A-D. Figure 4.1E shows a phase diagram demarcating the transitions between the different wetting regimes.

### v2

@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [189] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecule [190, 191, 192, 54], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [193] and porous networks [194]. The complex compositional features resulting from this mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different types of droplet within a cell.
@@ 3.1 Complex molecular interactions
Condensates contain a huge variety of components, including water, ions, small peptides, proteins and flexible polymers such as RNA and DNA [59, 22]. Since it is challenging to directly study interactions between these molecules in their native environment, much experimental work has thus far relied on in vitro reconstitution [195, 196, 126, 37]. Here, it is crucial to exploit label-free techniques [197, 198] since fluorescent labels typically strongly affect droplet condensation [198, 199]. One key insight from these approaches is that often a few scaffold components control the formation of condensates, while client molecules partition into them [200, 201, 202, 203, 204]. However, this dichotomy is likely an oversimplification and it will be interesting to more rigorously quantify the contributions of individual components, for example by using the recently developed dominance metric [205], which can uncover basic interaction motifs between components [206].
@@
It is clear that biomolecular phase separation involves many different kinds of interactions between molecules [207]. To distill the essential features of these interactions, it is often helpful to coarse-grain particular microscopic models [88, 208]. For instance, many biomolecules can be conceptualised as associative polymers [209, 210, 211, 212, 81, 207] and described by sticker-and-spacer models [209, 213, 82, 22, 214], where flexible spacers link stickers with specific interactions. However, there are also many other approaches to modelling the complex interplay of biomolecules that leads to phase separation [22, 85, 59, 215]. Such models allow us to elucidate how particular molecular features, like the amino acid sequence of proteins, affect condensate properties [216, 217, 218, 219], and can explain why dilute phases often exhibit surprisingly large clusters [220, 221].
@@
Interactions between biomolecules can be roughly divided into weak, transient interactions and strong, persistent bonds. While weak multivalent interactions are generally responsible for phase separation into liquid-like phases, strong interactions tend to form long-lived assemblies, which can persist as nano-domains inside condensates [222]. Moreover, chemical modifications of biomolecules can affect their interactions. For instance, proteins can be changed by post-translational modifications [189], which are implicated in condensate regulation [223, 53, 224, 35]. Such changes can induce multimerisation, which can in turn enable phase separation. To understand this phenomenon, we imagine a molecule whose weak interactions are insufficient for phase separation. A multimer of such molecules, however, can interact with many other multimers, which effectively increases the molecular volume $\Omega$ while maintaining the interaction strength $\chi$ per unit volume [165]. This process essentially reduces the entropic cost of concentrating molecules, promoting phase separation. This feature has been used to induce artificial condensate in cells [13, 225]. More generally, it is likely that combinations of weak and strong interactions enable fine-tuning of condensate properties [226, 227].
@@ 3.3 Complex interfaces
Interfaces between droplets and the surrounding fluid are characterised by a sharp change in composition over a finite spatial width. Typically, a rapid and continuous exchange of components between droplets and the surrounding fluid occurs across the interface, even at equilibrium. Since droplets in biological cells generally comprise long, complex molecules, however, this exchange may be limited by an interface resistance. Such a resistance could result from potential barriers, a reduction in diffusivity [260], saturation of binding sites [261], or electrostatic effects [262].
Many proteins molecules are amphiphilic, and can thus act as surfactants, preferentially associating to the interface [263, 264, 265] to form Pickering emulsions. In general, surfactants stabilise droplet interfaces, resisting both dissolution and coalescence. The formation of complex interfaces via preferential protein association can therefore improve structural integrity of condensates and assist in size regulation [266], as well as leading to complex droplet morphologies [267]. Note that although surfactants usually result in kinetic droplet stabilisation, some complex particles may also result in thermodynamic stabilisation [268, 269].
@@ 3.4 Electrostatic effects
Charge is another important biomolecular property that is involved in the regulation of condensates [274]. On the one hand, charge patterns on protein surfaces can mediate specific, short-ranged interactions [275]. On the other hand, a net charge induces electrostatic interactions that can potentially range over long distances. Indeed, many condensates contain charged components [276, 277], which affects interfacial properties [278, 279, 262, 280], slows down coarsening [281, 282], limits droplet size [155, 283, 284], influences droplet–environment interactions [285, 286], and can potentially lead to droplet division [287, 288]. The dynamics of charged species can be described by the classical electrostatic framework. The central quantity is the charge density, $\varrho(\boldsymbol{r})=\sum_{i}q_{i}\phi_{i}$, which depends on the volume fractions $\Phi=\bigl{(}\phi_{1},\ldots,\phi_{{N_{\mathrm{c}}}}\bigr{)}$ of all species $i=1,\ldots,{N_{\mathrm{c}}}$, each with charge $q_{i}$ per unit volume. The charge density $\varrho$ induces an electrostatic potential $\varphi(\boldsymbol{r})$, which satisfies the Poisson equation
@@
where $\varepsilon$ is the electric permittivity. The potential $\varphi$ then gives rise to an electric field $\boldsymbol{E}=-\nabla\varphi$ and modifies the overall free energy,
$F_{\mathrm{e}}[\Phi(\boldsymbol{r})]=F_{\mathrm{bulk}}[\Phi(\boldsymbol{r})]+\frac{1}{2}\int\varphi(\boldsymbol{r})\varrho(\boldsymbol{r})\text{d}V\;,$ (3.5)
where the first term is given by equation (2.6), while the second term captures electrostatic effects. The associated electrochemical potential, $\mu_{\mathrm{e}}=\nu\delta F_{\mathrm{e}}/\delta\phi_{i}$, gains a term proportional to $\varphi$, $\mu_{\mathrm{e}}=\mu+q_{i}\varphi$, relative to the uncharged case. This electrochemical potential shows that positively charged molecules move down gradients in the electrostatic potential $\varphi$. In particular, small mobile ions will tend to distribute so that the system becomes charge neutral, inhibiting long-ranged electrostatic effects. In the simplest case, these ions can be described by Debye-Hückel theory, which accounts for their translational entropies and electrostatic interactions [289, 290]. The main result is that electrostatic effects do not extend significantly beyond the Debye length $\lambda_{\mathrm{D}}=[\varepsilon k_{\mathrm{B}}T/(\sum_{i}c_{i}q_{i}^{2})]^{1/2}$, where $c_{i}$ is the mean number concentration of the ion species $i$ with charge density $q_{i}$. Note that ion concentrations can be phase-dependent [274], meaning that electrostatic interactions may be longer ranged than would be expected from the average salt concentration [284].
Over short distances, electrostatic effects mediate attraction between oppositely charged molecules, which can lead to phase separation from a neutral solvent via complex coacervation [291, 25, 292, 293]. Coacervates have been studied for about a century [294] and have been implicated in the origin of life [295, 296]. In the simplest case, the attraction of oppositely charged molecules can be captured by an additional term in the local free energy $f$ given by equation (2.1). For instance, in $d=3$ dimensions, Voorn–Overbeek theory leads to the free energy density [297, 298]
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of the electrostatic interaction. Here, $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$ is the Bjerrum length, which describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [299]. Since the permittivity $\varepsilon$ of cytosol is comparable to water [300], a typical Bjerrum length is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{nm}$. We thus have $\alpha\sim 0.1\ldots 1$ for typical molecular volumes of $\nu_{0}\sim 10^{3}\ldots 10^{5}\,\mathrm{nm^{3}}$, indicating that electrostatic effects can lower the free energy significantly.
@@ 3.5 Multiphase coexistence
Cells often comprise many different types of condensates, consistent with Gibbs phase rule, which limits the number of different phases to be smaller than or equal to the number of different components; see section 2.1.1. The number of coexisting phases and their compositions can generally be determined by minimising the free energy of the entire system, which leads to the coexistence conditions given by equations (2.5). However, even for the simple Flory–Huggins theory given by equation (2.1), these relations are difficult to solve since they are non-linear and depend on all interaction parameters $\chi_{ij}$ and compositions $\bar{\phi}_{i}$. It is thus not surprising that the resulting phase diagrams, showing the equilibrium phases as a function of the system composition, can be complex [103]. Even in the simplest case of two interacting solutes, the topology of the phase diagram depends on the interaction strength; see Figure 3.1. This multi-component, multi-phase problem has been tackled by linear stability analysis of the homogeneous state [301, 302, 303, 118, 304], grand-canonical simulations [305, 306], continuation methods [307], convexification of the free energy density [103], and machine learning methods [308].
@@ 3.5.1  Random mixtures
where $\alpha>0$ is a shape parameter. The equimolar case is recovered in the limit $\alpha\rightarrow\infty$; $\alpha=1$ samples all compositions uniformly; and $\alpha<1$ result in mixtures with a few dominating species. Taken together, this protocol enables study of the average behaviour of a multi-component liquid with random interactions.
@@ 3.5.2  Instability of homogeneous states
Phase separation of a multi-component mixture must occur when the homogeneous state is unstable, which is the case when the Hessian $H_{ij}=\partial^{2}f/\partial\phi_{i}\partial\phi_{k}$ exhibits at least one negative eigenvalue. If there are multiple negative eigenvalues, the eigenvectors generally point in different directions, suggesting that the system segregates into various phases. The number of negative eigenvalues, $N_{\mathrm{u}}$, can be predicted for the Flory–Huggins model, for which
$H_{ij}$ $=\frac{k_{\mathrm{B}}T}{\nu_{0}}\left(\frac{\delta_{ij}}{\Omega_{i}\phi_{i}}+\frac{1}{\Omega_{0}\phi_{0}}+\chi_{ij}\right)\;,$ (3.8)
following from equation (2.1). Using the random interactions and compositions introduced in section 3.5.1, random matrix theory can be used to predict $N_{\mathrm{u}}$ [309, 301, 118]. For a large number of components and uniformly distributed compositions ($\alpha=1$), the average value of $N_{\mathrm{u}}$ is approximately [310]
@@ 3.5.3  Coexisting equilibrium phases
The scaling predicted for the number of unstable modes and the phase count is consistent with the results for binary phase separation discussed in section 2.1.3: equation (2.18) implies that the fraction above which the homogeneous state becomes unstable scales as $\chi^{-1}$, while equation (2.21) shows that the minimal fraction for forming a droplet scales as $e^{-\chi}$. The same scalings are captured by equations (3.9) and (3.10), suggesting that this is a universal feature of phase-separating systems [310]. This scaling also implies that a stability analysis of the homogeneous state does not carry reliable information about the phase behaviour of multi-component liquids.
@@ 3.5.4  Towards realistic systems
The scaling result given in equation (3.10) predicts some features of condensate coexistence observed in cells, but is likely not quantitatively accurate. This discrepancy is because the cell’s composition is clearly not random, but instead homeostatically controlled, and interactions between biomolecules are also not random, but structured [312, 206, 313, 314, 315] and also under cellular control [316, 97, 317, 318, 319]. Moreover, interactions are typically complex [320], likely involving higher-order interactions [321] and the effect of polymer dynamics [322, 212, 210]. Note also that the number ${N_{\mathrm{p}}}$ of coexisting phases is only the most basic descripter of multi-component phase separation. The compositions of different phases are also crucial for cellular function since they give rise to the droplets’ identities. Along these lines, it will also be interesting to understand what determines the interfacial parameters $\kappa_{ij}$ in equation (2.6), since these are crucial for predicting surface tensions and thus multiphase morphologies [323, 103, 324]. In the simplest case, surface tension and other related quantities might simply scale as a power law of the distance to a critical point [325], but whether this situation holds generally is unclear. Besides surface tension, droplet morphology is also shaped by other objects in the surrounding environment, as discussed in section 4 below.
@@ 4.1.1  Partial, full and pre- wetting
Figure 4.1: Different phenomena resulting from droplet–structure interactions. (A) Dense phase partially wets the external boundary, forming a compact droplet whose interface meets the boundary at a contact angle $\vartheta$. The areas of the interfaces between the droplet and the dilute phase, and the droplet and the boundary are denoted by $A$ and $S$, respectively. (B) Dense phase completely wets the boundary. (C) In an undersaturated system, droplet material pre-wets the boundary, forming a layer of thickness $h$. (D) Droplet volume fraction $\phi$ as a function of distance from the boundary. Note that the gradient of $\phi$ at $z=0$ is the same for all three cases. (E) Phase diagram showing the pre-, full and partial wetting regimes. The typical size of the pre-wetting regime has been exaggerated for clarity. (F) Droplet–droplet interactions between two immiscible phases result in either droplet separation, engulfment, or partial wetting (clockwise from top left). Panels (A) to (E) inspired by figures in [327] and panel (F) inspired by figure in [13].
@@
In a supersaturated system ($\bar{\phi}>\phi^{(0)}_{\mathrm{out}}$), $\Delta\!f>0$, and so macroscopic droplets are able to form regardless of the presence of external boundaries. Nevertheless, surface interactions still have a significant impact on droplet regulation, both in terms of droplet nucleation (see section 4.1.4 below) and droplet morphology. In particular, if $\gamma_{\mathrm{in}}<\gamma_{\mathrm{out}}$ and $\gamma<\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}$, the dense phase will completely ‘wet’ the surface, forming a macroscopic layer between the surface and the dilute phase. This configuration is known as perfect, or full, wetting. For systems in which $\gamma>|\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}|$, the dense phase will instead partially wet the surface to form spatially compact droplets. The crossover between these two cases is known as the wetting transition [328]. The different phenomenologies of pre-, partial, and full wetting are highlighted in figures 4.1A-D. Figure 4.1E shows a phase diagram demarcating the transitions between the different wetting regimes.